# Effect of Adaptive Communication Support on Human-AI Collaboration

### v1

@@ Abstract
Effective human-AI collaboration requires agents to adopt their roles and levels of support based on human needs, task requirements, and complexity. Traditional human-AI teaming often relies on a pre-determined robot communication scheme, restricting teamwork adaptability in complex tasks. Leveraging strong communication capabilities of Large Language Models (LLMs), we propose a Human-Robot Teaming Framework with Multi-Modal Language feedback (HRT-ML), a framework designed to enhance human-robot interaction by adjusting the frequency and content of language-based feedback. HRT-ML framework includes two core modules: a Coordinator for high-level, low-frequency strategic guidance, and a Manager for task-specific, high-frequency instructions, enabling passive and active interactions with human teammates. To assess the impact of language feedback in collaborative scenarios, we conducted experiments in an enhanced Overcooked-AI game environment with varying levels of task complexity (easy, medium, hard) and feedback frequency (inactive, passive, active, superactive). Our results show that as task complexity increases relative to human capabilities, human teammates exhibited stronger preference towards robotic agents that can offer frequent, proactive support. However, when task complexities exceed the LLM’s capacity, noisy and inaccurate feedback from superactive agents can instead hinder team performance, as it requires human teammates to increase their effort to interpret and respond to the large amount of communications, with limited performance return. Our results offers a general principle for robotic agents to dynamically adjust their levels and frequencies of communications to work seamlessly with human and achieve improved teaming performance.
Introduction
@@
Human-robot collaboration has been extensively studied and applied across diverse scenarios, demonstrating strong potential towards enhanced efficiency and performance (???????). As task complexity increases, agent adaptability becomes increasingly essential for seamless teamwork. Previous work has developed methods and tools for robot to adapt their actions based on inferred human objectives (?), trust level (?), and individual preferences (?). Recent work also begin to incorporate language-based feedback (??) to enable more direct communications and lower user barriers. However, robot communications in these approaches primarily focused on relative simple commands such as “pick up the book” or “move left a little”, which do not reflect the level of human-robot communications required in real-world applications.
@@
Recent advancements in large language models (LLMs) have brought powerful reasoning (????), natural language understanding (??), contextual awareness (?), and generalization capabilities (?), enabling more advanced, prolonged communications (???). These methods has empowered agents to process ambiguous and complex instructions from human (?), engage in more natural and dynamic conversations (??), and learn from a diverse set of inputs (??).
@@
However, even in these LLM-enhanced communications, human teammates continue to play a predominant role in requesting specific tasks and providing suggestions during collaboration (?). Schoenegger et al. (?)’s study suggested that state-of-the-art LLMs can often match or surpass human performance in various domains. Based on these results, we hypothesize that allowing LLM agents to more proactively participate in (?) or even initiate communications with human teammates can enhance teaming performance and efficiency.
@@
To test this hypothesis and systematically evaluate the impact of different forms of language feedback provided by robots on collaboration efficiency and human satisfaction, in this study we develop HRT-ML, a human-robot teaming framework incorporating multi-modal language feedback to support dynamic, context-aware teaming styles. To enable the robot to provide effective language feedback, HRT-ML comprises two main modules: a Coordinator, which manages overall collaboration strategies and delivers low-frequency or passive instructions and feedback, and a Manager, which determines appropriate subtasks based on the coordinated plan at each stage, offering high-frequency instructions. Combining Coordinator and Manager allows the robotic agent to provide different forms of instruction at different frequencies. To investigate how the form and frequency of language feedback influence teaming performance, we performed user studies using four different agent active levels: Inactive, Passive, Active, and Superactive. We find that as the environment becomes more challenging, participants exhibit stronger preference for agents that provide active support, whereas in simpler tasks, frequent agent feedback is often perceived negatively, reducing overall performance and human satisfaction. These findings reveal the importance for the active level of robot language support to dynamically adapt based on task complexity and team capability. Inaccurate or overly-frequent feedback can decrease team efficiency, as participants must take extra time to understand and correct these suggestions.
@@
  • performed user studies to determine how forms and frequency of agent language feedback influence teaming performance
  • discovered how the active level of robot language support should adapt based on task complexity and team capability, to best enhance teaming performance
To the best of the authors’ knowledge, this is the first study that systematically explores the effect of different types of LLM-based multimodal language feedback on teaming performance across a wide range of task complexities.
Testbed: Overcooked-AI
To test the influence of language feedback on human-robot collaboration, we chose Overcooked-AI (?), a platform designed to assess multi-agent coordination skills. In this game, players are motivated to collaborate actively to maximize their score by completing orders within a time limit. A score of 60 points was awarded when the correct soup is served. Partial points will be awarded to incomplete or incorrect soups based on the number of missing/incorrect components. To cook a soup, chefs need to finish specific subtasks in sequence according to the recipes (see Fig. 1A), and to finish subtasks, chefs need to move and interact with the environment. This process can be applied to any collaborative setting: first, reasoning through subtasks to achieve the overall goal, then selecting low-level atomic actions to complete each subtask.
Figure 1: (A) Cooking process to complete an order. (B) The designed human-AI collaboration interface (left: game layout, right: communication panel). The red cross represent an example of the intermediate empty counter used to collaborate.
Subtasks
The subtasks in an overcooked environment for a single agent can be broken down into three main parts:
@@
    Cooking: Once the ingredients are placed in the pot, the agent has to start cooking. A timer on the pot signals when the soup is ready to be served.
@@
In multi-agent collaborative scenarios, each agent has a different path cost for completing each subtask, such as picking up an onion versus a tomato. Some subtasks may be unachievable for certain agents. Furthermore, by decomposing a subtask into multiple smaller subtasks that can be performed by multiple agents, the overall time cost can be reduced. For example, in Fig. 1, the green agent might place the onion at the red cross point (5,2), allowing the blue agent to pick it up from there and add it to the pot.
Atomic Action
@@
To finish a specific subtask, such as picking up an onion, the agent must execute a sequence of atomic actions, including movement commands like up, down, left, right, stay, and interact for picking up or placing objects.
Figure 2: Flowing Human-Robot Teaming Framework (HRT-ML). It contains two modules: the Coordinator and the Manager. The Coordinator considers human preferences, formulates overall strategies, and provides low-frequency feedback to guide the collaboration. The Manager processes these strategies along with the state information to generate high-frequency subtask instructions for both human players and greedy planners, facilitating efficient task coordination and execution.
@@
Human-Robot Teaming Framework with Multi-Modal Language Feedback
To provide adaptive language feedback in human-robot collaboration, HRT-ML includes two core components: the Coordinator and the Manager. The Coordinator leads high-level strategy discussions with humans, generating a final coordination plan and offering low-frequency feedback. In contrast, the Manager handles detailed subtask allocation and provides high-frequency feedback to guide human players. Humans and other agents will receive the feedback and execute low-level actions.
Coordinator
The Coordinator is responsible for designing overall collaboration strategies and discussing them with the human partner. To support this, we employed a structured chain-of-thought approach to make the suggestions and discussions more effective. It begins by retrieving relevant information, including the layout, the overall collaboration goal, rules, and the agent’s state formatted as a text description (Figure 2) using a prompt template, and then queries GPT-4o with it. It first analyzes the possible subtasks required to achieve the overall goal, then evaluates the difficulty of each subtask for each agent, and generates a plan that maximizes collaboration efficiency. For example, based on the environment (Fig. 1B), our Coordinator suggests: Given the layout, the human (blue agent) should focus on picking up the tomato and serving the dish, while I will handle the onion and place the dish at location (5, 2) or (6, 2). We note that the coordination querying is made conversational, allowing humans to continuously revise the plan through continuous discussion after reviewing the plan. During these coordination discussions, the game remains paused until the human chooses to end the conversation. The conversation can be initiated proactively by the Coordinator at a low frequency or by the human. When a human request is made, we incorporate their suggestions into the coordination plan prompt for initially querying GPT-4, explicitly asking it to evaluate the feasibility of these preferences and propose a corresponding coordination plan.
Manager
The Manager assigns subtasks at each stage, providing high-frequency support to guide the human toward the final goal. First, the Manager uses a subtask filter to identify feasible subtasks based on the agent’s and environment’s current states, selecting from the potential subtask list from the overall plan created by the Coordinator. Next, it converts the selected subtasks, coordination plan, current agent states, and environment state into a language-based prompt, querying GPT-4 to generate the next subtask. This query is generated immediately upon completing the current subtask, with an average latency of $\sim$ 1.2 seconds. The determined subtask in every query will then be converted into language feedback and sent to humans.
Greedy Planner
Once given the target subtask, the human determines the atomic actions needed to complete it. For collaborative tasks with an autonomous agent, a greedy planner using Depth First Search (DFS) finds the optimal path with the lowest action cost to finish the subtask.
Multi-Modal feedbacks
@@
Building on the proposed HRT-ML framework, we introduce agents that provide four different language feedbacks described as follows: Inactive Feedback, Passive Feedback, Active Feedback, and Superactive Feedback.
@@
  • Passive Feedback agent (PFA): The PFA starts to provide passive feedback only when a human requests. The human player takes the role of the leader, while the agent acts as the follower, passively responding to human requests. If there are no specific human commands, PFA will behave like an IFA.
  • Active Feedback agent (AFA): The AFA collaborates with humans as peers. In this mode, both humans and agents can reach out to give language feedback. For user study, we prompt the GPT-4o coordinator in low frequency ($\sim$ 20s) to analyze human conversation history and suggest coordination strategies.
@@
In this section, we aim to explore the agent’s ability to provide language feedback to help improve human satisfaction and teaming efficiency. Based on the proposed HRT-ML, we consider four types of language feedback and three different layouts: easy, medium, and difficult, with increasing task difficulty and map complexity (Fig. 3). Human participants will play on each layout paired with an autonomous agent with varying levels of language feedback. Further details will be provided in the following section.
Layouts with different complexities
@@
To test the performance of the four agents and the influence of language feedback, we implemented four overcooked maps (Fig. 3). One of the maps is an introductory map, designed for participants to get familiar with the game and operation. The other three maps have varying levels of difficulty, which we refer to as easy, medium, and hard. The more challenging maps have more “dead-ends”, requiring humans and agents to have a better coordination strategy to maintain the team efficiency. Furthermore, the hard map introduces complexity in task orders by requiring multiple ingredients, which demands that humans and the agent reason about orders containing both tomatoes and onions.
@@
Figure 3: Overview of the human study procedure involving 16 participants, beginning with an introduction map to learn game mechanics and agent behavior followed by trials on Easy, Medium, and Hard Layout with four different agents. Post-session surveys were conducted to collect data on participant satisfaction, engagement, trust, and feedback.
Participants

### v2

# Effect of Adaptive Communication Support on LLM-powered Human-Robot Collaboration

@@ Abstract
Effective human-robot collaboration requires robot to adopt their roles and levels of support based on human needs, task requirements, and complexity. Traditional human-robot teaming often relies on a pre-determined robot communication scheme, restricting teamwork adaptability in complex tasks. Leveraging strong communication capabilities of Large Language Models (LLMs), we propose a Human-Robot Teaming Framework with Multi-Modal Language feedback (HRT-ML), a framework designed to enhance human-robot interaction by adjusting the frequency and content of language-based feedback. HRT-ML framework includes two core modules: a Coordinator for high-level, low-frequency strategic guidance, and a Manager for subtask-specific, high-frequency instructions, enabling passive and active interactions with human teammates. To assess the impact of language feedback in collaborative scenarios, we conducted experiments in an enhanced Overcooked environment with varying levels of task complexity (easy, medium, hard) and feedback frequency (inactive, passive, active, superactive). Our results show that as task complexity increases relative to human capabilities, human teammates exhibited a stronger preference towards robotic agents that can offer frequent, proactive support. However, when task complexities exceed the LLM’s capacity, noisy and inaccurate feedback from superactive robotic agents can instead hinder team performance, as it requires human teammates to increase their effort to interpret and respond to a large number of communications, with limited performance return. Our results offer a general principle for robotic agents to dynamically adjust their levels and frequencies of communications to work seamlessly with humans and achieve improved teaming performance.
[…]
1 Introduction
@@
Human-robot collaboration has been extensively studied and applied across diverse scenarios, demonstrating strong potential towards enhanced efficiency and performance [20, 10, 29, 15, 35, 24, 25]. As task complexity increases, robot adaptability becomes increasingly essential for seamless teamwork. Previous work has developed methods and tools for robot to adapt their actions based on inferred human objectives [25], trust level [8], and individual preferences [3]. Recent work also begin to incorporate language-based feedback [28, 31] to enable more direct communications and lower user barriers. However, robot communications in these approaches primarily focused on relative simple commands such as “pick up the book” or “move left a little”, which do not reflect the level of human-robot communications required in real-world applications.
@@
Recent advancements in large language models (LLMs) have brought powerful reasoning [36, 38, 1, 16], natural language understanding [23, 34], contextual awareness [12], and generalization capabilities [14], enabling more advanced, prolonged communications [4, 27, 18]. These methods has empowered robots to process ambiguous and complex instructions from human [23], engage in more natural and dynamic conversations [34, 19], and learn from a diverse set of inputs [14, 32].
@@
However, even in these LLM-enhanced communications, human teammates continue to play a predominant role in requesting specific tasks and providing suggestions during collaboration [23].  Schoenegger et al. [30]’s study suggested that state-of-the-art LLMs can often match or surpass human performance in various domains. Based on these results, we hypothesize that allowing LLM-powered robots to participate more proactively in [33] or even initiate communications with human teammates can enhance teaming performance and efficiency.
@@
To test this hypothesis and systematically evaluate the impact of different forms of language feedback provided by robots on collaboration efficiency and human satisfaction, in this study we develop HRT-ML, a human-robot teaming framework incorporating multi-modal language feedback to support dynamic, context-aware teaming styles. To enable the robot to provide effective language feedback, HRT-ML comprises two main modules: a Coordinator, which manages overall collaboration strategies and delivers low-frequency or passive instructions and feedback, and a Manager, which determines appropriate subtasks based on the coordinated plan at each stage, offering high-frequency instructions. Combining Coordinator and Manager allows the robotic agent to provide different forms of instruction at different frequencies. To investigate how the form and frequency of language feedback influence teaming performance, we performed user studies using four different robot active levels: Inactive, Passive, Active, and Superactive. We find that as the environment becomes more challenging, participants exhibit a stronger preference for robots that provide active support, whereas in simpler tasks, frequent feedback is often perceived negatively, reducing overall performance and human satisfaction. These findings reveal the importance for the active level of robot language support to dynamically adapt based on task complexity and team capability. Inaccurate or overly frequent feedback can decrease team efficiency, as participants must take extra time to understand and correct these suggestions.
@@
  • performed user studies to determine how forms and frequency of robot language feedback influence teaming performance.
  • discovered how the active level of robot language support should adapt based on task complexity and team capability, to best enhance teaming performance.
To the best of the authors’ knowledge, this is the first study that systematically explores the effect of different types of LLM-based multimodal language feedback on teaming performance across different task complexities.
2 Related Work
Robots serving as teammates and collaborating with humans in shared workspaces is an emerging paradigm in human-robot interaction [9, 13, 39]. Recent advancements in LLMs have further enabled robots to interpret human intent, generate task plans, and engage in more natural interactions [37, 22, 39]. In our work, we specifically focus on the role of communication in LLM-based human-robot collaboration, as effective information exchange is crucial for task coordination, intent alignment, and adaptive interaction.
Communication in human-robot collaboration can generally be categorized as either unidirectional or bidirectional. In unidirectional communication, robots do not proactively engage with humans; instead, they infer human intent from instructions and make autonomous decisions based on observed cues [36, 6, 22, 23]. In contrast, bidirectional communication involves robots actively interacting with humans—seeking clarification, sharing updates, and negotiating task execution strategies [16, 37, 26, 39]. Several studies have explored unidirectional and bidirectional communication strategies, finding that humans perceive robots capable of bidirectional communication as more intelligent and collaborative [2, 39].
However, existing studies [2, 39] on bidirectional communication primarily focus on updating status or requesting simple, predefined subtasks without incorporating complex reasoning about the coordination strategies. A well-defined bidirectional communication framework for complex coordination tasks is still lacking. Moreover, the impact of different forms and frequencies of bidirectional communication on coordination performance and human preference remains unexplored.
3 Testbed: Overcooked-Environment
To test the influence of language feedback on human-robot collaboration, we chose Overcooked-AI [5], an environment designed to assess multi-agent coordination skills. In this game (Figure 1, humans (blue hat agents) and robots (green hat agents) are motivated to collaborate actively to maximize their score by completing orders within a time limit. A score of 53 points was awarded when the correct soup was served. Partial points will be awarded to incomplete or incorrect soups based on the number of missing/incorrect components. To cook a soup, chefs need to finish specific subtasks in sequence according to the recipes (see Fig. 1A), and to finish subtasks, chefs need to move and interact with the environment. This decision making process can be applied to any collaborative setting: first, reasoning through subtasks to achieve the overall goal, then selecting low-level atomic actions to complete each subtask.
Figure 1: (A) Cooking process to complete an order. (B) The designed human-AI collaboration interface (left: game layout, right: communication panel). The Blue hat agent and Green hat agent are operated by humans and robots, respectively. The red cross represents an example of the intermediate empty counter used to collaborate.
3.0.1 Subtasks
The subtasks in an overcooked environment can be broken down into three main parts:
@@ 3.0.1 Subtasks
    Cooking: Once the ingredients are placed in the pot, the chef has to start cooking. A timer on the pot signals when the soup is ready to be served.
@@
In collaborative scenarios, the human and robot have a different path cost for completing each subtask, such as picking up an onion versus a tomato. Some subtasks may be unachievable for humans or robots. Furthermore, by decomposing a subtask into multiple smaller subtasks that can be performed by human and robot together, the overall time cost can be reduced. For example, in Fig. 1, the robot might place the onion at the red cross point (5,2), allowing human to pick it up from there and add it to the pot.
3.0.2 Atomic Action
@@
To finish a specific subtask, such as picking up an onion, the chef must execute a sequence of atomic actions, including movement commands like up, down, left, right, stay, and interact for picking up or placing objects.
Figure 2: Human-Robot Teaming Framework with Multi-Modal Feedback(HRT-ML). It contains two modules: the Coordinator and the Manager. The Coordinator designs overall coordination strategies using Directed Acyclic Graph (DAG) and dynamically add and delete nodes and edges through communication with humans. Three types of coordination—graph structure revision, graph node attribute revision, and adding temporary subtasks—are enclosed in a red dashed rectangle. A blue subtask node represents tasks that are ready for execution, a yellow node indicates tasks that are not yet ready, and a red node signifies emergency tasks. The action cost between two subtasks is indicated with numbers on the edges. The Manager interprets the graph-represented coordination strategy to allocate subtasks to humans and robots. It provides subtask-level instructions to guide humans toward the overall goal and updates the graph status upon the completion of each subtask.
@@
4 Human-Robot Teaming Framework with Multi-Modal Language Feedback
To provide adaptive language feedback in human-robot collaboration, HRT-ML leverages a Directed Acyclic Graph (DAG) to represent subtasks and its dependencies (e.g., finishing an onion soup) for human-robot collaboration. Based on the DAG, HRT-ML includes two core modules: the Coordinator and the Manager. The Coordinator handles high-level strategy discussions, generating an initial subtask graph and revising it based on conversation. The Manager is responsible for allocating detailed subtasks based on the finalized coordination DAG, monitoring the status of assigned subtasks for both the human and the robot, and updating the graph accordingly. Once a subtask is determined, a Greedy Action Planner is used to generate atomic actions for the robot.
4.1 Subtask Graph for Human-Robot Collaboration
We represent the collaboration goal as a dynamic directed acyclic graph (DAG) (Figure 2) where nodes represent subtasks (e.g., pick onion, cook onion), and edges represent dependencies between tasks. Each edge is defined as (parent node, node, cost), where the cost is calculated using the Greedy Action Planner (Section 4.4). Each node in the graph contains attributes for different coordination purposes:
  • Task id: Each task gets a unique id, such that the LLM only needs to return the task id, and we can retrieve all other information based on the task id.
  • Task Type: Categorized as Putting, Getting, or Operating. There exist constraints when assigning different subtasks to robots or humans depending on their state, e.g., a robot that already held an onion can not execute any Getting or Operating tasks.
  • Status: Represents the execution state, includes ready-to-execute, executing, success, failure, emergency. The status of each subtask is considered when an LLM is prompted to assign subtasks for Humans and Robots.
  • Target Position: Specifies a list of locations in the environment where the task can be performed. One subtask can be executed at multiple locations. This will enable humans and robots to explore additional collaboration strategies, such as the human placing an onion on an empty counter while the robot retrieves it.
  • Notes: The subtask node includes a notes attribute that captures human preferences or additional instructions (e.g., the human chef prefers to perform picking up an onion; the human chef is observed to be proficient at picking up onion.).
  • Parent subtasks: The parent subtasks indicate the prerequisite subtasks that are required to be success before this subtask can be executed.
  • Priority: Priority is considered when assigning subtasks to humans and robots. It is calculated as the accumulated cost from the current node to the sink (termination) node.
  • Running time: The running time acts as a counter, recording the duration for which the robot has been assigned to this subtask.
The subtask graph enables the robot to translate language-based coordination into modifications within the graph, helping the LLM organize conversations and align its knowledge with the human’s.
4.2 Coordinator
The Coordinator of the robot is responsible for designing overall collaboration strategies and discussing them with the human partner. The initial subtask graph is generated using a one-shot LLM prompt (Appendix 12.1), incorporating recipe and kitchen layout information. We then implement a conversation-based graph modification, which can be initiated by the robot at a low frequency with the corresponding prompt (Appendix 12.1) (active mode / superactive mode described in Section 4.5 or proactively by the human through language-based feedback). The coordination process can be divided into the following decision tree:
  • The robot first identifies the type of human request, which falls into one of three categories (Figure 2, red dashed rectangle); if it is unsure, it will further ask a human to clarify.
    • Subtask Graph Structure Change: The human expresses thoughts on how to complete the overall task. This may involve adding nodes, further decomposing a subtask for human and robot execution, or correcting incorrect graph dependencies. For example, based on the environment (Figure 1B), one of the possible coordination strategies is that the robot can place the onion at location (5, 2), and the human (blue hat) should focus on picking up onions at (5, 2).
    • Subtask Attribute Change: The human specifies preferences for subtask assignments. For example, based on the environment (Figure 1B), a human says that he would like to handle picking up onions and dishes as well as putting them on the counter (5, 2) and (6, 2).
    • Temporary Subtask Assignment: The human is not designing an overall collaboration strategy but instead requests the robot to execute a specific temporary subtask. For example, if the robot is blocking the human’s path while they attempt to put an onion in the pot, the human may instruct the robot to move to a specific position.
  • After identifying the type of coordination, the subtask graph is updated using the prompts shown in Appendix 12.1 for each category.
During these coordination discussions, the game remains paused until the human concludes the conversation.
4.3 Manager
The Manager (Figure 2) is responsible for real-time subtask allocation based on the information from the subtask graph and subtask status monitoring.
4.3.1 Subtask allocation
The Manager were first asked to analyze the current state of humans and robots along with the list of subtasks that are ready to execute in the subtask graph. The prompt (Appendix 12.2) explicitly instructs the robot to consider task priorities and types based on the current state, as well as human-reported preferences for each subtask. In addition to subtask assignments, the Manager is also prompted to generate language instructions to guide the human to execute their assigned subtasks toward the final goal. This feedback is selectively presented to the human based on the feedback mode described in Section 4.5.
4.3.2 Subtask status update
The Manager will be triggered to evaluate the current executing task to update the status given the human and robot action trajectory when an Interact action is executed (Prompt details can be found in Appendix 12.2). Once a subtask is identified as success, the subtask nodes without unresolved dependencies will be marked as ready-to-execute, while others remain not-ready until their prerequisites are completed.
4.4 Greedy Action Planner
Once a subtask (target location) is assigned, the robot uses a greedy planner based on Depth First Search (DFS) to determine the sequence of atomic actions required to complete the task. The planner computes the optimal path with the lowest action cost, considering the current environment layout and chef states.
4.5 Multi-Modal Feedbacks
@@
Building on the proposed HRT-ML framework, we introduce robotic agents that provide four different language feedbacks described as follows: Inactive Feedback, Passive Feedback, Active Feedback, and Superactive Feedback.
@@
  • Passive Feedback Agent (PFA): The PFA starts to provide passive feedback only when a human requests. The human player takes the role of the leader, while the robotic agent acts as the follower, passively responding to human requests. If there are no specific human commands, PFA will behave like an IFA.
  • Active Feedback Agent (AFA): The AFA collaborates with humans as peers. In this mode, both humans and robots can reach out to give language feedback. For user study, we prompt the GPT-4o coordinator in low frequency ($\sim$ 20s) to analyze human conversation history and suggest coordination strategies.
@@ 5 Data collection
In this section, we aim to explore the robot’s ability to provide language feedback to help improve human satisfaction and teaming efficiency. Based on the proposed HRT-ML, we consider four types of language feedback and three different layouts: easy, medium, and hard, with increasing task difficulty and map complexity (Fig. 3). Human participants will play on each layout paired with a robotic agent with varying levels of language feedback. Further details will be provided in the following section.
5.1 Layouts with Different Complexities
@@
To test the performance of the four robotic agents and the influence of language feedback, we implemented four overcooked maps (Fig. 3). One of the maps is an introductory map, designed for participants to get familiar with the game and operation. The other three maps have varying levels of difficulty, which we refer to as easy, medium, and hard. The more challenging maps have more “dead-ends”, requiring humans and robots to have a better coordination strategy to maintain team efficiency. Furthermore, the hard map introduces complexity in task orders by requiring multiple ingredients, which demands that humans and the robot reason about orders containing both tomatoes and onions.
@@
Figure 3: Overview of the human study procedure involving 16 participants, the study begins with a tutorial map to introduce game mechanics. Followed by the tutorial, the human players will then play on the easy, medium, and hard layouts with four agents with different language feedback. Post-session surveys were conducted to collect data on participant satisfaction, engagement, trust, and feedback.
5.2 Participants